Good morning, I have a problem that I need help with, thanks so much.
Jose has savings accounts for each of his five children.
They contain $225, $126, $328, $75, and $25. Find the median and the mean amount of money per account.

Answer by checkley77(12844)

Arrange these figures from smallest to largest.
25,75,126,225,328.
The middle term is the median or the $126 account.
adding all these accounts & then dividing the toptal by 5 gives the mean.
(25+75+126+225+328)/5=779/5=155.8 is the mean amount.